Conquest of Elysium II

Johan Karlsson e-mail d92jk@efd.lth.se

970827

Contents

1 Installation 1

2 Playing the game 1

3 Modes of play 1

4 Game instructions 2

5 Tax 2

6 Character classes 2

6.1 Warlords . . . . . . . . . . . . . . . . 3 6.1.1 Baron . . . . . . . . . . . . . 3 6.1.2 Barbarian . . . . . . . . . . . 3 6.1.3 Maharaja . . . . . . . . . . . 3 6.1.4 Senator . . . . . . . . . . . . 3

6.2 Magic Users . . . . . . . . . . . . . 4 6.2.1 Necromancer . . . . . . . . . 4 6.2.2 Demonologist . . . . . . . . 4 6.2.3 Warlock . . . . . . . . . . . 5 6.2.4 Enchanter . . . . . . . . . . 5

6.3 Priests . . . . . . . . . . . . . . . . . 5 6.3.1 Cardinal of El . . . . . . . . 6 6.3.2 High Priestess of Baal …. 6 6.3.3 Witch . . . . . . . . . . . . . 6 6.3.4 Shaman . . . . . . . . . . . . 7 6.3.5 Druid . . . . . . . . . . . . . 7

6.4 Non Humans . . . . . . . . . . . . . 7 6.4.1 Troll King . . . . . . . . . . 7 6.4.2 Orc warlord . . . . . . . . . 7 6.4.3 Dragon Lord . . . . . . . . . 8 6.4.4 Dwarf King . . . . . . . . . . 8 6.4.5 Elf Queen . . . . . . . . . . 8 6.4.6 Elf King . . . . . . . . . . . 8

7 Monsters 8 7.1 Spells . . . . . . . . . . . . . . . . . 8 7.2 Fear . . . . . . . . . . . . . . . . . . 8 7.3 Charm . . . . . . . . . . . . . . . . 8 7.4 Curse . . . . . . . . . . . . . . . . . 9 7.5 Siege engines . . . . . . . . . . . . . 9 7.6 Charge . . . . . . . . . . . . . . . . 9 7.7 Energy drain . . . . . . . . . . . . . 9 7.8 Immortality . . . . . . . . . . . . . 9 7.9 Resistances . . . . . . . . . . . . . . 9 7.10 Magic Items . . . . . . . . . . . . . 10

8 How to win 10

1 Installation

You need a colour monitor or tv, an ATARI Falcon and at least 4 megs of memory to run this program. If you have got these things the installation will be very simple. All you have to do is copy all files into a single directory of your hard drive and double click on coe.prg to run the game.

2 Playing the game

Each player controls a powerful warlord or wizard whose objective is to take control over the continent of Elysium. The game can be played by one to eight players and if there are less than eight players the computer can control some or all of the remaining players.

3 Modes of play

After starting the game you will have to choose which game-mode to play. The buttons are Random world, Customized world, Scenario, Load game and Quit. You may press the buttons by using the mouse or the keys 1-5. If you choose to customize your world you will have to choose the size of the world and what kind of society there shall be. There are six different societies to choose from.

The first society is the dark ages when man came to Elysium. Ruins of long dead civilizations lay scattered through the realm and dragons roam the lands. Elven courts and dwarven cities are not yet destroyed. Human influence is weak and the forces of nature strong. There are few cities in these early times.

The second society is the agricultural times when man inherited the earth. Cities are rare, but there lies hamlets and villages in every corner of the lands of Elysium. It is a time of heroes and pioneers.

The third society is the empire. The empire was formed after the hard times of the pioneers. A great city was buildt and others founded to fill the need of the growing population. A new kind of military unit was trained to quench the uprisings of the realm and to fight the barbarians who poured down from the north. The villages of the empire has grown into cities and the forests are being cut down.

The fourth society is the interregnum. No empire can last for ever. Battlefields and ruins of the empire litter the lands. Few cities are left and not too many villages either. The human lands are in disorder and only a powerful leader will be able to rebuild the empire.

The fifth society is the monarchy. In towers of stone sits feudal lords gathering powers for their king. There are lots of fortresses and towers in the monarchy, but the cities of the empire has yet to be rebuildt.

The sixth society is the dawn of the new empire. Cities are once again growing and the future of humankind looks bright and promising. The hard years are forgotten and the leaders of the lands are hopeful.

4 Game instructions

When you have selected what kind of game you want to play and what players are participiating you are ready to begin the struggle. First you will see your home citadel in the middle of the screen, surrounded by lots of black unknown terrain. In your citadel you will have one or more commanders, select one by clicking on the square or by clicking on the commander button. After that press the button with crossed swords on to select which men are to acompany this commander. Each commander has three action points to spend by walking, spell casting etc. You may walk to any adjacent square not containg water by clicking on it while the commander is selected.

If you press the right button on a square you will see all people that are in that square. If you click on a figure you will get information about the figure chosen. The information revealed is melee, attacks if more than one, hitpoints, magic resistance, movement mode and health. In some cases you will also see some pictures at the bottom of the screen. These pictures are items or resistances (see below). If you double click with your right mouse button you will get information on the terrain in that square (or you may single click if there is no army in that square). All locations that are owned by someone will have a coloured shield in the top left corner and all locations containing an army will have a coloured sword in it. There are two other symbols that might appear in a square - there is a cross when the square is blessed and a small temple if a temple has been constructed there.

These are some keyboard shortcuts for commonly used functions.

• q Quit turn

• s Recruit men

• t Change tax

• c Change followers

• p Use special power

• o Use this location’s special power

• n Next commander

• d Remove commanding ability (when viewing his stats)

• g Give item away (when viewing item)

• 1-5 Change game speed

• s (in fight) speed up fight

• m Enter menu

• Esc Quit game and return to desktop

5 Tax

The tax determines what your monthly income will be. If you tax your subjects heavily for a long period, the people will eventually revolt. The peoples mood are not only affected by income tax, winter makes them sour, spring makes them happy and rulers using them as sacrifices makes them very angry indeed.

6 Character classes

There are nineteen different charcter classes you may choose from, everyone has different strengths and weaknesses.

6.1 Warlords

When mortal men entered the Elysium it was a perilious realm. Powerful sorcerers who had gathered knowledge over centuries lived in ruined castles and monsters roamed the forests. Even the animals were large and hostile to mankind. To protect the newly arrived pioneers some men built fortresses of stone and timber. These men became leaders of their kin and was called warlords. The warlords of different regions of Elysium developed different tactics and unique troops to guard their lands. The warlords however could not use the powers of magic and the sorcerers became the greatest threat to their newly founded realms.

6.1.1 Baron

The human border vas defended by landowners known as barons. Some barons were warlike and restless. As humankind humankind spread over Elysium and became dominant they tried to usurp their masters powers and waged war on their neighbours. The large well trained forces at their disposal made this possible and soon they were in control of large parts of the human territories. But not all humans were content with this share. The sorcerers of Elysium gathered armies of their own and nonhuman warlords called out for vengeance. The barons must once again try to subjugate Elysium.

This warlord has a very wide variety of human troops and huge siege engines. The knights are formidable units who uses the charge to reduce the enemy ranks in the initial phase of the battle. They are however vulnerable to spearmen, and these are common in cities and villages. Heavy infanterists are strong and do not occupy two spaces as do the knight, nor do they need to fear spearmen. Elites are ordinary troopers trained by a weapon master. The elites are effective against hordes of weak creatures due to their great number of attacks. The baron knows how to keep his subjects in line and can whip peasants into obedience. His realm is seldom plagued by revolts and mine strikes. He also knows the value of a milita in defending the lands of the peasants.

Abilities: The peoples of the barons lands will not be as upset as others under heavy tax. The baron and the high lords can build siege towers and wooden towers when standing in a forest. In wooden towers you can buy troops. The baron and the high lords can also raise levies in villages and towns. The more inhabitants the settlement has the more levies you get.

6.1.2 Barbarian

Time after time barbarian hordes has plundered the lands of Elysium. Their sudden arrival at the fringes of civilazation sends tremors through the realm. Where they comes from no one knows, some says from small villages in secluded areas, others say from the netherworlds. The barbarian threat is always present and only with the forces of the barons can the tide be turned.

The barbarians are physically more powerful than ordinary men, but they are also more vulerable to magic. The barbarian warlord cannot buy advanced weapons like siege engines but may acuire minotaurs and werebears to strengthen their troops. The werebears can hide in ordinary troops and can surprise unwary opponents. The barbarian troops are cheaper than ordinary humans and you will probably be leading large armies.

Abilities: None.

6.1.3 Maharaja

The barons of the north learned how to train soldiers and build war engines. This made them sucessful in the art of war. But while the barons competed for dominance of the riches in the north they failed to se what lies hidden in the southern jungles. Here maharajas formed small kingdoms hidden in the jungles and deserts. By domesticating wild beasts and using them in their armies they were able to dominate the south. The fearsome war elephants are the most powerful of all regular troops and in combination with archers they are deadly. Somtimes the maharajas ally with rakshasas, tiger headed lords of the jungle.

The maharaja uses war elephants and tigers in warfare. The elephants travel fast in the jungles and deserts and will not die on you unless you are outmatched. Other special troops are the snake charmer who summons snakes to do his battles and the fakir who uses a rope to entangle opponents. The rakshasa, finally, is a powerful magician who do not require much gold, but will consume your elephants, one a year. Abilities: None.

6.1.4 Senator

During the time of the empire a new kind of tactics evolved. Troops were designed to fight humans dissidents and to eliminate the barbarian threat. The soldiers were called legionaires and they were joined together in tight formations called cohorts.

In the cohort the newly recruited legionaires were placed next to veterans to learn the art of war. The system was very effective against the humans they were supposed to fight, but there were other forces in Elysium.

The cohort is a formidable unit by human standards. The tight formation gives the cohort unmatched number of attacks and a good capacity to take physical punishment. The senator sometimes has the opportunity to buy gladiators or net wielding retarii. Besides the ordinary wizards available to every warlord the senator can hire augurs, priest diviners of the empire. The well diciplined legionaires are capable builders and can construct siege engines, towers and fortresses where there is wood enough.

Abilities: The senator is immobile. The centurions can construct siegetowers and outposts like the baron. The outpost however, is a citadel. If all citadels are lost you are out of the game but thanks to the outposts the senator is more likely to survive. The augur can use augury to gain information from distant places. The location viewed is random.

6.2 Magic Users

Long ago Elysium was a wild land untouched by human hands. Elves and trolls fought each others and dwarves buildt halls of splendour under the mountains of the land. Wild animals roamed the forests and even the plants were sentinent. To this realm came a magician of great renown, the Archmage called, and opened a gate to Terra, the land of the humans. Through this gate human sorcerers came with their strange magic and they settled in the rich lands of Elysium. As the ages went by these magicians grew in power and knowledge and when the gate to Terra suddenly widened and men came forth, the sorcerers used the magic and resources of Elysium to twart the warlords threatening the stability of the realm.

6.2.1 Necromancer

The Necromancers are sorcerers adept in the dark art of necromancy, the summoning and creation of living dead. Only humans can be forced to take this cursed resemblance of life. Thus necromancers were unheard of until the coming of man. The Necromancer may raise the buried dead or the fallen in battle to strengthen his army. This practice is forbidden by the gods and the necromancers who are foolish enough to disturb the dead will loose their sanity in the process. The other aspect of necromancy is the harvest of hands. In stormy nights darkly robed men collects the left hands of hanged murderers. These are the hands of glory. The burnt flesh of such hands is pleasant to ghosts and evil spirits who can aid the necromancer in his horrible art. The hands can also be used in even worse ways. When enough of the hands are burnt spirits of the darkest realms can be called upon to instruct the necromancer how to transform his body into that of a living dead.

The necromancer has very weak troops in the beginning of the game. It is essential to the necromancer to find a battlefield or graveyard early in the game. The dark knowledge is important in this search. The animated dead are weak and will never heal. This makes them unrealible and not to useful. The flesh heap on the other hand is the strongest animated monster in the necromancers army. It will not heal but absorb the flesh of enemies in close combat. In animating the dead the apprentice is nessesary, otherwise you will go mad and loose the ability to take rational action. The summoned undead requires hands of glory which can be found in cities, villages and gallows. The ultimate goal of most necromancers is to transform into a powerful undead and live forever!

Abilities: The necromancer has four major powers. With ”raise dead” you animate some of the dead in the square you are standing. Humans who has died in battle adds to the numbers of dead in the square. Every attempt to raise dead leads you further from sanity and so does every dead raised. If you raise lots of dead you lose more sanity. If there are lots of dead in a place you will probably get yourself a mummy or perhaps a flesh heap. These are more powerful animated dead and replace a bunch of lesser undead. When you get insane you will temporarily lose control of your actions. The only means to get rid of insanity is to become undead yourself. The second power is the summoning of more powerful undead. This costs hands of glory. The major undead are often commanders. The third power is dark knowledge. This spell costs a few hands of glory and it lets you find a place with lots of dead. The fourth and last power is the transmutation of yourself into a powerful immortal undead. This act restores your sanity. Some spirit infested places will allow you to summon undead and transform yourself at lower costs.

6.2.2 Demonologist

The Demonologist is a mage priest who summons the inhabitants of abyss and inferno. The spirits of these worlds are powerful and horrible to behold. They crave human flesh and can only be bound into service if they are sated. The greatest of these demons are unsatiable and cannot be bound unless the demonologist is lucky indeed and the sacrifice is enormous. If the demonologist fails the binding ritual, the demon will attack and try to consume the weak human who tried to rob him of his freedom.

The forces at the demonologists disposal are truly powerful. Lesser demons are physically very strong and some of them has magical powers as strong as any magic user. Greater demons are even more powerful and the mighty demon lords are practically unequalled in magic abilities as well as in physical combat. Many demons use fire to damage their opponents and fire resistant troops can be a problem to the unwary demonologist.

Abilities: There is three classes of demons to choose from. The more powerful the demon is the more human sacrifices they crave. You can choose to sacrifice small, normal and huge amounts of humans. The more you use the greater the chance to succeed in controlling the demon.

6.2.3 Warlock

The Warlocks were among the first humans who arrived in Elysium. They found that elemental power is trapped in the gems of Elysium. By freeing this power the warlock can create or attract elemental spirits to serve him. Rubies contain fiery power, diamonds aerial energy, saphires watery power and emeralds earthly force. When large amounts of gems are used powerful elemental spirits will arrive. The warlock is also known to bind elemental force to his own body, giving him powers beyond that of ordinary men.

The elementalist has some very strong beings at his service. The elemental lords (efreet, marid, djinn and dao) are nearly as powerful as a demon lord. They are commanders and may summon elementals. Giants and elemental beings are strong but often lacking in magical powers. Elementals are comparable with lesser demons.

Abilities: The warlock collects gems of different types. Gems can be found in mines and in cities. The gems lets them use five different powers. Create elemental lets you create one larger or several smaller elementals. Larger elementals can be bound to your body. This destroys the elemental. If you bind air you gain the power of flight, earth gives you hitpoints, fire grants you the power of fire in battles and water soothes wounded humans. The greater beings cost more gems and are more powerful. They cannot be bound.

6.2.4 Enchanter

The Enchanter has learned the art of making constructs. The constructs are mago-mechanic, living statues, armors or swords. The constructs are huge and can withstand enourmous amounts of damage. With the help of these mechanical giants the enchanter can crush even the mightiest of armies. But these titans are extremly expensive. To build them you need whole mines of minerals. The lesser constructs are less expensive and more commonly used when building armies.

The enchanter will have troops with enormous amounts of hitpoints. As the constructs cost money to create, the enchanter often lacks in ordinary troops, but even a few golems can crush the mightiest armies. Constructs do not heal, but has lots of resistances.

Abilities: The enchanter can build constructs. This costs money. Some constructs requires special reources (see below). If a mine is used to build a golem, the mine is emptied of all minerals, a lake will be dried and a marsh will remain, and so on. The construction and enchanting of objects is a very time craving work and costs several action points. The requirements for construction are these.

• necrotod, dead people

• flesh golem, dead people

• tree golem, a forest

• oak golem, an ancient forest

• clay golem, a marsh or swamp

• ice golem, a fresh water lake

• stone golem, a coal mine

• iron golem, an iron mine

• crystal golem, a silver mine

• onyx golem, a gold stream

• gold golem, a gold mine

When constructing a golem the enchanter has to be in the required terrain or, in the case of a lake, next to it.

6.3 Priests

The priests are religious leaders of man. They have the ability to communicate with their gods and to interpret the signs of the gods. The priests strive to gain control over the lands to please their gods.

The priests of El are the most common in the times of the empire and the new age. The priestesses of Baal are common in the interregnum and in the monarchy. The three priests connected to nature are most common in the early times.

6.3.1 Cardinal of El

Before man there were few monotheistic religions. With man came new faiths claiming the one truth. The worship of El was new in Terra and the adherents of the faith were hunted down and killed. In Elysium the faith flourished. Whith the rise of the empire the religion became the most withspread and even the emperor was converted. With the strength of the adherents an inquisition was formed to eradicate differing faiths. The primary target of the inquisition was the worshipers of Baal, the other monotheistic faith of Terra. But with the fall of the empire the faith was weakened. If El could not help resisting the destructive forces of Elysium, why then worship him. People turned instead to the old religion of Terra, the worship of Baal.

The templars are strong units by human standards. They charges and has a good number of hitpoints. Many of the cardinals troops has increased magic resistance. Ordinary units can be blessed. If your congregation is large enough you may be granted an angel. Angels are powerful beings with commanding abilities and priestly powers. The inquisition and the four riders are fearsome but out of your control. They travel from citadel to citadel to punish the infidels with the wrath of El.

Abilities: The cardinal of El collects adherents to his faith. The more congregation he has got the more powers are bestowed upon him by El. When you have a small congregation you can only buy templars. With growing congregation you can bless areas to rais the income there by 0.5 gold. With even larger flocks you will be able to heal the wounded, bless soldiers and build temples from which you gain temple taxes. Temples can only be buildt in population centres. If your congregation is large enough El will grant you a holy warrior to fight for your cause. With large congregations you might even get a inquisition. When your congregation is ready you will have the possibility to call down the four horsemen of the apocalypse to destroy the infidels and faithless. This costs you your life. The archbishops and bishops has got the same powers as the cardinal but you will need more congregation for their abilities to work.

6.3.2 High Priestess of Baal

In old times in Terra the god Baal incarnated and let the humans taste fear and destruction. The moon went blood red and snow fell for three years. To placate the horrible god men started to sacrifice their fellow men. This pleased the god who consumed the souls of the sacrificial victims. Baal withdrew into the world from which he came and fed upon slaughtered souls. Since this time mortal men has worshiped and sacrificed to the horror god. Baal has let his prescence be known from time to time. When the Martyr of El called out to the poor and famished, Baal sent his subjects to war. The children of El were hunted down and killed and the followers fled into the promised land of Elysium. The Baalite faith was weak in Elysium and it was only with the fall of the empire, as the faith established itself. During the interregnum the Baalite faith became strong. Now, in the dawn of the new empire it seems, El has again taken up the challenge for supremacy.

The priestess of Baal are among the most powerful characters. The hideouds creatures summoned at the higher levels are nearly as powerful as demon lords and they need not be controlled. The lower beings are not so strong but may arrive in great numbers.

Abilities: The priestess of Baal collects sacrifices. By sacrificing victims she is rewarded by Baal, the hungry god. Horrible monsters with feeding rage are the most common reward. These monsters need not be controlled as they act on the will of their father and master, Baal. The priestess can also curse the very lands of Elysium. Wither wood destroys the forest you are standing in. Call winter changes the season to winter. Fimbul winter is a three year winter and armagheddon reanimates all buried dead in the world. Invoke Baal lets your god incarnate in your body. With this comes a fimbul winter for free.

6.3.3 Witch

In the ages long before the arrival of men there were beings of tremendous power roaming Elysium. These beings were unchallenged lords of the yet unnamed realm. In time they tired of the lands and melded into the earth to sleep. They slept for ages and other beings came to the realm. The sounds and laughter of these new inhabitants disturbed the sleeping ones and sometimes one of them awoke to punish the merry ones. The sleeping ones became feared and worshiped by the people of Elysium. The adherents of this old faith are called witches and they are feared by elves and man alike. The witches collects fungii in the forests and marshes of Elysium. With these they brew a decoct pleasant to the Old Ones. The smell of the cauldron makes them hungry and reminds their slow minds of the times before, when all was quiet and the land was theirs. In rage and hunger the Old Ones and the beings close to them follows the witch to reclaim their lands from the noisy and the quick.

The beings at the witches disposal are physically strong, but lacking in ranged attacks. Archers or crossbowmen are crucial. The monsters are not very difficult to get. You will also be aware of your surroundings as you collect forests.

Abilities: Witches uses fungii collected in forests and marshes. During autumn the amount is increased and during winter you gain nothing. When you summons one of your hideous beings you will have to choose the amount of fungus to use, the less fungus you uses the easier it will be to loose control over the summoned creatures.

6.3.4 Shaman

The Shaman is a tribal priest who guides the tribe by asking the spirits for advice. By leaving his body he can travel through the realm og the spirits and confront spirits of evil and good. While in the spirit world he can see distant places and gain knowledge of ages gone. From this other realm the shaman can possess living beings and replace their souls with a spirit under his command. By sending nightmares, evil dream spirits, to a victim, he can reduce the bond between soul and body enabling him to install a false soul. The false soul totally controls the actions of its host being. Shamans are known to lead barbarians, but they lack the leadership to collect armies.

The shaman is never really strong in brute force. On the other hand he can attack from afar. The totem pole has a strong magic attack, but is immobile and guards the home. Greater spirits are very strong but they are alone more often than not. The collection of herbs increases your awareness as you owns forests.

Abilities: The shaman collects herbs in forests. They gain increased numbers during summer and none during winter. The herbs can be used to scry, send nightmares, possessions and to summon spirits. The ability to scry lets you know distant places. When scrying choose one square to scry upon and you will see what is there as if you owned that place. The nightmares are only used to prepare a victim for possessions. The possession lets you take control of one victim with limited magic resistance. If you succed you gain control of the being. If it is a commander it retains its abilities as commander. The spirit attacks will send spirits to destroy enemy troops. This can be useful when you are about to attack an army. The spirit disappears after combat.

6.3.5 Druid

The druid is a priest of nature and a caretaker of the lands. As the lords of Elysium found cities the druids calls the animals of the realm to do battle on the exploiters. The animals of Elysium are strong and has souls like men. The druid aids the animals and make descisions for their sake. By collecting herbs and using them in strange rituals the druid can call the beasts to his side. Ultimately he can gain acceptance of the powerful beings of the legends or the guardians of the forests.

Some animals are stronger than humans, but they lack ranged attacs. The legendary beings are powerful as are the guardians of the forests.

Abilities: The druid uses herbs like the shaman. Which animals are summoned depends on the terrain he summons in, e.g. in the plains there are lots of rabbits and in the jungle tigers may be found. If he has a huge amount of herbs he may attempt to summon legendary beings such as rocs and basilisks in the mountains or guardians of the forests in the ancient forests. The druid can also heal dead forests.

6.4 Non Humans

Ages ago, long before the coming of man, there were other races inhabiting the lands of Elysium.

6.4.1 Troll King

The Troll King is the physically most powerful warlord. He is capable of con- quering towns on his own. He can also throw large boulders at city wall just like a catapult. His regeneration enables him to heal at an extraordinary rate, just like his main companions, the trolls.

6.4.2 Orc warlord

The orcs are very strong fighters, however they are also more easily affected by magic than humans. The orc scouts are expert forest burners and may destroy forests to stop people from getting herbs and fungus there.

6.4.3 Dragon Lord

The Dragons are among the oldest inhabitants of Elysium. Where they came from is unknown, but their prescence is seldom so. Dragons are greedy beasts capable of capturing dwarwen cities by themselves. When in control of the dwarven hoard they lies upon it and sleeps for ages. The dragons are thus not useful in warfare. The dragonspawn or the draconians are quite the opposite. The draconians are erect reptilian humanoids with leathery wings and tough scaly bodies. They vary greatly in colour and physical power depending on their ancestral lines. The green draconians are the most common and they are the base unit in dragonspawn armies. The whites are the lowest ranking of all and used mostly as workers. The blue are very strong and works in the mines and in places where strength is needed. The red are the warrior caste and fearsome in battle, they are somewhat rare. The noble caste is parted in two, the black warrior nobles and the purple nobles schooled in the arts of magic. The royal golden draconians call themselves dragon lords and they know how to summon their ancestors, the true dragons of times gone. The draconians are semi nomadic and are not unlike the barbarians sweeping down from the north from time to time. The draconians however tend to sweep in over a mine, summon one of their ancestral dragons and fly on to another mine.

The draconians are strong troops with lots of attacks. They all has some resistance, green against poison, white against frost, and so on. All draconians can fly. This means you can fly over mountains without loosing more action points than one. The dragons summoned to guard the mines are among the most powerful creatures in the game. But even a dragon will be hard pressed if attacked by an army with creatures resistant to their breath. The draconians are rare troops and you will not always be able to buy one. The black and purple nobles are rare. If you want one of these you should hoard your money for some time. White and green are the most common.

Abilities: The Dragon Lord may summon dragons to guard his mines. The dragons takes a part of the income in return for their service. This summoning costs money. With richer mines you get stronger dragons.

6.4.4 Dwarf King

The dwarves are exellent miners and the Dwarf King get higher income from mines than the other warlords.

6.4.5 Elf Queen

The Elf Queen collects life force from forests and helpful creatures will appear automatically within her lands (the more life force, the more powerful creatures). The life force varies according to the seasons.

6.4.6 Elf King

The Elf King collects life force from forests and helpful creatures will appear automatically within her lands (the more life force, the more powerful creatures). The life force varies according to the seasons.

7 Monsters

There are more than three hundred different monsters in this game, so it will probably take a year or two before you have seen them all! Many of the monsters have special abilities (there are over one hundred different special abilities!), here is an exlanation of some of them.

7.1 Spells

Most magic-users are able to cast spells in combat. Spell casting is very difficult if the caster is in meele combat and lesser magic users won’t be able to cast any spells from the front rank. Many spells can be resisted and spells are easier to resist the more magic resistance you have got and the lousier the spell caster is.

7.2 Fear

This magic attack scares one or more opponents if they fail a magic resistance check. If a commander is scared away, he will loose his ability to function as a commander permanently. This also applies to sleep and stun attacks. The Ghost has an extremely powerful fear attack that is impossible to resist, however it is possible to survive it and only flee in panic if the victim doesn’t die from fear. The victims that have been scared will survive if their companions win the battle.

7.3 Charm

This is an attack against the opponents magic resistance. Should the attack succeed, the victim will become charmed and join the attacker. A commander that is charmed will loose their ability to command.

7.4 Curse

This is a spell used by some evil religous sects. It turns the victim into a zombie, ready to do its masters bidding. The spell only works on humans, any non-human affected by the spell will simply die if they get affected.

7.5 Siege engines

Some siege engines may be used to destroy the walls of a fortress without en- gaging the defenders. When a commander has got one of these machines in his army he gets a catapult button. By pressing this button he may bombard any adjacent square containing a fortress.

7.6 Charge

Knights and cavalry uses a powerful charge to engage the enemy. This charge is only used in the first round of combat. The drawback to this attack is that some spearmen sets has developed a technique to make the charge less favorable. By setting their spears into the ground when the charge come, the knights will charge into these and take some damage.

7.7 Energy drain

Some of the more powerful undeads e.g. Wights and Vampires drain energy from the victim it touches. The energy drained will be used to heal the undead, should it be wounded. A more powerful energy drain called incorporate, takes the victims energy and uses it to grow in strength. By incorporating victims the creature may become stronger than he initially was.

7.8 Immortality

Some powerful beings e.g. Vampires and Liches are immortal. When killed in combat they will immediately return to their original home citadel (or their owners citadel if they haven’t got one of their own) and after some rest they will be as new. If the home citadel has been conquered by an opponent, the immortal can be permanently slayed.

7.9 Resistances

Some creatures possess strong endurance against the elements or against certain magical attacks due to strong personality or lack of soul. Undead creatures for example cannot be frightened. These resistances (or weaknesses) are shown at the bottom of the info screen of a certain being. The resistances are as follows:

• Armour. Reduces damage in combat by one or more steps. Symbol: one or more

shields.

• Poison resistance. The creature is immune to poison damage.

Symbol: a green droplet.

• Fire resistance. The creature is immune to fire damage. Symbol: a fire.

• Cold resistance. The creature is immune to frost damage. Symbol: icicles.

• Shock resistance. The creature is immune to electricity damage. Symbol:

a lightning bolt.

• Immolation. The creature is surrounded by a fiery aura and each attacker

recieves fire damage. Symbol: a large fire with a figure in it.

• Aura of cold. The creature is surrounded by a icy wind and each attacker

recieves frost damage. Symbol: a snowy whirl.

• Vulnerability to fire. The creature takes double damage from fire attacks.

Symbol: a burning stick.

• Vulnerability to cold. The creature takes double damage from frost attacks.

Symbol: an ice covered stick.

• Divine healing. The creature heals at increased rate. Symbol: a bloody bandage.

• Regeneration. The creature regenerates in combat and heals all wounds after battle.

Symbol: a hydra head.

• Never heals. The creature cannot heal. Symbol: a red droplet.

• Immortality. The creature will return to your home citadel if killed. Symbol:

a phoenix immolating.

• Fear immunity. The creature cannot be frightened by any means. Symbol: an

anguished face.

• Wakefulness. The creature cannot fall asleep by any means. Symbol: a bed.

• Charm resistance. The creature cannot be charmed by any means. Symbol: a

broken heart.

• Undead. The creature can be turned by priests of El. The undead also

has the following resistances: Fear immunity, Charm resistance, Wakefulness and Poison resistance. Symbol: a skull.

7.10 Magic Items

Hidden in old ruins and in the care of great magicians or brave heroes magic items can be found. These items grant magical powers and can strenghten the lucky owner.

When you find an item you has the opportunity to choose a reciever. As some items are cursed and impossible to trade you should be careful when choosing reciever. The item will then be shown above resistances on the info screen. If you click on the item you will get some information about it. Many items grant resistances and powers in combat. Others increase your hitpoints or your melee value. When looking at an item you can give it away by hitting the ”g” key.

8 How to win

You are out from the game when either all your commanders are killed or when you don’t have a citadel any longer. To win you just have to be the last player left.